Circular No. 6933

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SGR 1900+14 B. R. Oppenheimer, J. S. Bloom, S. S. Eikenberry, and K. Matthews,

California Institute of Technology, communicate: “We have obtained narrowband 2.2-μm images of the proposed infrared counterpart to SGR 1900+14 (Vrba et al. 1996, Ap.J. 468, 225) on June 3.486 UT, 4.1 days after the recent bursting activity reported on IAUC 6929. Differential photometry of stars ‘A’ and ‘B’ (the double M stars) and star ‘C’ of Vrba et al. yields m_{A} − m_{B} = +0.4 ± 0.1 and m_{A} − m_{C} = +4.5 ± 0.2 mag, consistent with the measurements of Vrba et al. Over the course of 10 min (ten separate frames), no variability was detected in the double M stars to a limit of 0.1 mag.”

XTE J2012+381 P. Callanan and J. McCarthy, University College, Cork; and M. Garcia

and J. McClintock, Harvard-Smithsonian Center for Astrophysics, report: “Observations of the XTE J2012+381 error region (IAUC 6920, 6927) were performed using the Cerro Tololo 1.5-m telescope (+ CIRIM detector) on June 1.36, 2.35, 3.35, and 4.36 UT. The J and K magnitudes of the object near the radio position (IAUC 6924) were 15.0 ± 0.1 and 14.3 ± 0.1, respectively. No variability at more than the 10-percent level was observed in the J and K bands between each night. Closer inspection of those images with a FWHM of < 1′′ reveals an apparent southeast-northwest elongation of this object, indicating that this is a composite image of two stars ∼ 1′′ apart. This confirms the detection of the faint red companion to USNO 1275.13846761 found by Hynes and Roche (IAUC 6932). As the position of the northwest star is consistent with the improved radio position (IAUC 6932), it may be the true counterpart to the x-ray source. Further observations, especially in conditions of good seeing, are encouraged.”

XTE J1748–288 B. A. Harmon, M. L. McCollough, C. A. Wilson, S. N. Zhang, and

W. S. Paciesas, Marshall Space Flight Center, report for the BATSE Team: “This x-ray transient is being detected with BATSE, with the following average daily fluxes in the bands 20–30 and 50–70 keV: June 3, 150 ± 65 and 160 ± 45 mCrab; June 4, 450 ± 70 and 280 ± 45 mCrab. There was no detection on June 2 (< 75 mCrab). The spectrum extends to at least 100 keV and softened significantly between June 3 (power-law photon number index −2.2 ± 0.3) and June 4 (index −2.9 ± 0.1).”

1998 June 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green